30 likes | 107 Views
Update on cryostat cavity operations from 8/07/2014 to 08/07/2014: Addressed cavity cavity issues, resolved power meter and polarity problems, identified PLL box fix, encountered multipacting effects, worked on temperature sensor data, and planned calibrations during cool-down. Late shifts due to fixing interlock system.
E N D
UK4R Status Maria Navarro, Karim Hernandez 8/07/2014
Update up to 08/07/2014 10:58 • Cavity in the cryostat at 4.15K • Few problems solved: • Power meter for the PT and PR were swapped in the switch box • AB’s had wrong polarity. *PLL box need to be fixed • Yesterday at 4.45K: • A lot of multipacting effects at very low field level ( ~ 0.1 MV), which did not allow us to get even the calibration point • We let the cavity for conditioning 3 hr before the scan • Scan of QvsU in pulsed mode • Now temperature sensor in the cavity are getting data but the sensors from cryo were dead. • Calibration for the cold part at the HOM and LOM frequencies. • Tobias also got the explanation for the double-exponential like decay that was observed in previous measurements
Update up to 08/07/2014 10:58 • At 8:30 start the cool down to 2.25 K • We are going to get calibration point during the cool down at : • different temperatures • different Frequencies • Problem with the interlock system. We are looking now to get this fixed • We need to work late, so we need to decide the time shifts